# Dual topological insulator with mirror-symmetry-protected helical edge states

### v1

@@ IV Topology analysis
Due to TRS, the total Chern number must be zero, i.e., $C=\left(C^{\uparrow}+C^{\downarrow}\right)=0$. By calculating the Chern numbers numerically, we obtain $C^{\uparrow}=-2$ and $C^{\downarrow}=2$ for our system, consistent with the second-order anti-skyrmion configuration in Fig. 3-c). This is related to the presence of the $Gk_{\pm}^{2}$ terms in the off-diagonal matrix elements of the $\bm{k}\cdot\bm{p}$ Hamiltonian (15). For $G=0$ (with the remaining parameters unchanged), the Chern numbers are given by $C^{\uparrow}=1$ and $C^{\downarrow}=-1$.
Another important topological invariant used to characterize 2D TIs is the so-called spin Chern number [25, 26, 27, 28, 90, 91, 92, 93], defined as
@@
For our system, $C_{s}=-2$, which characterizes the presence of a TI phase.
@@
As shown above, Na2CdSn hosts the TI and TCI phases simultaneously, thus it can be classified as a DTI protected by TRS and mirror symmetry. Note that both topological invariants (spin and mirror Chern numbers) are equal to $-2$ (in sharp contrast with the spin and mirror Chern numbers $-1$ reported in Ref. [70]). According to the bulk-boundary correspondence [23, 24], the Na2CdSn tri-atomic layer in a finite geometry must have two pairs of counter-propagating helical edge states at each of its boundaries [83]. We explore this feature in the next section.
@@ V Topological edge states
Let us now investigate the existence of topological edge states predicted by the non-trivial topological invariants calculated in Sec. IV. In Subsec. V.1, we analytically obtain the edge states for a semi-infinite plane geometry. In Subsec. V.2, we perform a numerical study of the topological edge states in a ribbon geometry.
@@ V.1 Semi-infinite plane
In this geometry, any edge-localized wave function $\Psi_{k_{y}}(x,y)$ must satisfy the following boundary conditions
@@
Equation (32) is the Born-von-Karman boundary condition, with $L_{y}$ the length of the system in the $y$ direction.
Consider the Schrödinger equation for the spin-up block, $h(k_{x}\rightarrow-i\partial_{x},k_{y})\psi(x,y)=E\psi(x,y)$, where $h(\bm{k})$ is defined by Eq. (25) and $\psi(x,y)$ is a particular two-component envelope function solution. The wave function $\psi(x,y)$ can be separated into a product of two factors, one corresponding to plane waves along the $y$ direction and the other corresponding to a non-periodic function along $x$. This allows us to write the following ansatz [96]
$\psi_{\kappa_{x},k_{y}}(x,y)=\frac{e^{ik_{y}y}}{\sqrt{L_{y}}}\begin{pmatrix}1\\ R_{\kappa_{x},k_{y}}\end{pmatrix}e^{i\kappa_{x}x},$ (33)
@@
Substituting (33) into the Schrödinger equation, we obtain four modes given by $\kappa_{x}=(-1)^{j}\kappa_{\pm}$, where $j=\{0,1\}$ and
@@
where $d_{k_{y}\pm}=-M-\bar{E}+(B-D)(\kappa_{\pm}^{2}+k_{y}^{2})$. From now on, we use $R^{j}_{k_{y},\pm}\equiv R^{j}_{\pm}$. Using Eqs. (33), (34) and (38), we write the general solution as the linear combination of the four modes
$\Psi_{k_{y}}(x,y)=\frac{e^{ik_{y}y}}{\sqrt{L_{y}}}\sum_{\begin{subarray}{c}j=0,1\\ l=+,-\end{subarray}}c_{j,l}\begin{pmatrix}1\\ R^{j}_{l}\end{pmatrix}e^{i(-1)^{j}\kappa_{l}x}.$ (39)
@@
Note that in applying the boundary conditions (30)-(32), we have already neglected the exponential terms that do not result in square-integrable solutions, i.e., that do not yield a vector Hilbert space. Using Eqs. (39)-(42), we obtain
$\Psi_{k_{y}}(x,y)=c_{0,+}\frac{e^{ik_{y}y}}{\sqrt{L_{y}}}\begin{bmatrix}e^{i\kappa_{+}x}-e^{i\kappa_{-}x}\\ R^{0}_{+}\left(e^{i\kappa_{+}x}-e^{i\kappa_{-}x}\right)\end{bmatrix},$ (43)
where $c_{0,+}$ is determined by normalizing the wave function. Similarly, for $\text{Im}\left(\kappa_{-}\right)<0$ we have
@@
$\Psi_{k_{y}}(x,y)=c_{0,+}\frac{e^{ik_{y}y}}{\sqrt{L_{y}}}\begin{bmatrix}e^{i\kappa_{+}x}-e^{-i\kappa_{-}x}\\ R^{0}_{+}\left(e^{i\kappa_{+}x}-e^{-i\kappa_{-}x}\right)\end{bmatrix}.$ (47)
@@
In Fig. 4, we show the analytical dispersion relations of the spin-up (red curve) and -down (blue) edge states [Eq. (48)] for the fitted parameters (Subsec. III.3). The gray-shaded areas represent the continua of delocalized bulk states (bulk region), whose boundaries are indicated by solid black lines.
Figure 4: (Color online) Topological edge states hosted by the $\bm{k}\cdot\bm{p}$ model in a semi-infinite plane geometry, with parameters from the fitting for Na2CdSn. The red (blue) solid lines represent the analytical solutions for the dispersion relations of the spin-up (-down) edge states [Eq. (48)], while the solid black lines delimit the bulk region, corresponding to the continua of delocalized bulk states (gray-shaded areas).
For each energy level inside the bulk band gap, there are two pairs of counter-propagating topological edge states of opposite spin. This is in agreement with the bulk-boundary correspondence theorem, considering the topological invariants calculated in Sec. IV. As it can be seen from equation (51), the quadratic dispersion relation of the topological edge states is a consequence of the non-zero parameter $G$.
Note that all edge states merge into the bulk region at $k_{y}\approx 0$. As $k_{y}$ moves away from zero, however, the energy of the edge states increases (in absolute value) slower than the energy of the bulk bands. Hence, here, the edge states never merge into the bulk region again. We emphasize that the effective model derived in Sec. III is only valid in the region close to $k_{y}=0$, thus we should not expect it to capture any realistic merging points that might occur at larger values of $k_{y}$. The absence of merging points is also observed when we consider edge states arising from the BHZ Hamiltonian in the absence of $k^{2}$ diagonal terms [97].
@@ V.2 Ribbon geometry
In Fig. 5-a), we show the electronic band structure for a ribbon of width $W=100\,$nm, i.e., a wide ribbon (compared to the lattice parameter $a=4.978\,$Å). Each $(k_{y},E)$-point is colored according to its spin, with red (blue) representing spin-up (-down) states. The energy bands forming a Mexican-hat-like structure constitute the bulk region. The energy gap $E_{g}\approx 285\,$meV between the lowest bulk-like conduction and topmost bulk-like valence bands is indicated in the figure.
@@
Figure 5: (Color online) Electronic states for Na2CdSn in a ribbon geometry. a) Electronic band structure showing spin-up (red) and spin-down (blue) states with a small spin-splitting caused by the bulk inversion asymmetry (BIA) and $\bm{k}$-dependent SOC. The bulk region is determined by the Mexican-hat-like energy bands, with a bulk band gap of $E_{g}\approx 285$ meV. The quadratic energy bands correspond to quasi-degenerate topological edge states, with the spin-split states localized at opposite sides of the ribbon. The colored cross marks indicate the selected edge states for our wave function analysis and the two black arrows denote the points of edge state crossings. b) Probability densities of the selected topological edge states as functions of $x$ (coordinate along the transversal direction). The color of the curves corresponds to the colors of the respective cross marks, while the solid (dashed) style represents spin-up (spin-down) states. c) Electronic band structure for Na2CdSn in a ribbon geometry of width $W=15\,$nm. The black arrows indicate the opening of an anti-crossing hybridization gap due to the finite-size effect.
@@
The quadratic energy bands crossing the bulk band gap correspond to the topological edge states. All energy bands are quasi-degenerate, including the edge states for which the spin-splittings are $\sim 10^{-1}$ meV. Similar to the spin-splitting of the energy bands in the infinite bulk system [Fig. 2-c)], the spin-splitting of the edge states is not accompanied by a mixing of the spin components due to the block-diagonal structure of the effective Hamiltonian [Eq. (15)].
@@
The quasi-degenerate topological edge states are localized at opposite sides of the strip. Note that, in the wide-ribbon limit, the energy dispersions of the edge states localized on the left (right) side can be mapped onto those of a vacuum/DTI (DTI/vacuum) semi-infinite system. In analogy to Subsec. V.1, we employ the envelope function approach and solve the problem for the DTI/vacuum geometry. We use the analytical solutions for the two semi-infinite systems to derive an expression for the inter-edge splitting. Subtracting the dispersion relations for spin-up and spin-down states, we show that the spin-splitting is given by $2b_{1,1}k_{y}$ and $2b_{2,1}k_{y}$ for the concave up and down bands, respectively.
@@
Notice that there is no structural inversion asymmetry in this case, which for the BHZ model is known to generate a Rashba term [99, 100, 101, 101, 102, 103] that lifts the degeneracy of the topological edge states [103]. For the ribbon geometry considered here, the spin-splitting arises solely from the BIA. It can be shown that the inclusion of BIA in the BHZ Hamiltonian [96, 104] also leads to inter-edge spin-splitting in 2D TI’s, although in this case the splitting is accompanied by a mixing of the spin components.
@@
In Fig. 5-b), we present the probability density of the edge states, indicated by the green and orange cross marks in 5-a), as functions of $x$ ($x=0$ and $x=W$ correspond to the ribbon extremities). Each cross mark represents two quasi-degenerate states localized at opposite sides of the ribbon (the line colors match the colors of the corresponding cross mark). The dashed (solid) green line corresponds to a spin-down (spin-up) state localized at the left (right) side of the ribbon. Similarly, the solid (dashed) orange line corresponds to a spin-up (spin-down) state localized at the left (right) side of the ribbon. Note that the states shown by the green color are slightly more localized than the ones represented in orange  [^2] A zoom-in on the probability density distributions shown in Fig. 5-b) reveals that the green (orange) curves present a normal (oscillatory) exponential decay with $x$. The decaying behavior of a given edge state is dictated by the region where it is found in the electronic band structure. This effect is well-discussed for the BHZ model in Ref. [107]. In our model, however, such an analysis is much more complicated. Therefore, we plan to provide a more detailed investigation of this topic in a forthcoming work..
@@
So far, we have only investigated the edge states for negative crystal momentum ($k_{y}\approx-0.24\,$nm-1 and $k_{y}\approx-0.48\,$nm-1). Since our model is invariant under TRS, each topological edge state has a Kramers partner localized on the same side of the ribbon, but with opposite momentum and spin. The yellow (black) cross mark in Fig. 5-a) indicates the two Kramers partners of the two states indicated by the green (orange) cross mark. We have verified that, in the wide-ribbon limit, the dispersion relations for the edge states localized at the left side of the ribbon agree exactly with those obtained for the semi-infinite plane geometry [see Eq. (48) and Fig. 4].
@@
For any energy level within the bulk band gap (e.g., at the Fermi level), there are four topological edge states at each side of the ribbon. On the left side, there are two spin-down (-up) edge states propagating along the $+y$ ($-y$) direction, while on the right side, there are two spin-up (-down) edge states propagating along the $+y$ ($-y$) direction.
The black arrows in Fig. 5-a) point to the two edge state crossings taking place around the same energy level, $E\approx 0.123\,$eV. At each crossing point, there are four topological edge states. Let us focus on the edge state crossing around $k_{y}\approx-0.4\,$ nm-1 (left black arrow). Two of the four edge states live on the same side of the strip, have opposite spins, and propagate along opposite directions. Since they live in different spin subspaces, they are orthogonal to each other and do not hybridize. The counter-propagating states localized at opposite sides of the ribbon, on the other hand, belong to the same spin subspace, and therefore, can hybridize. Naturally, the hybridization requires a superposition of their wave functions, the so-called finite-size effect [106, 107]. In Fig. 5-c), we show the electronic band structure for Na2CdSn in a ribbon of width $W=15\,$nm. For this narrow ribbon, the edge states (with the same spin) living on opposite sides of the strip hybridize. The hybridization gap due to the anti-crossing of the edge state bands around $k_{y}\approx-0.4\,$nm-1 and $k_{y}\approx 0.4$ nm-1 are indicated by the two black arrows in the figure.

### v2

@@ IV Topology analysis
Due to TRS, the total Chern number must be zero, i.e., $C=\left(C^{\uparrow}+C^{\downarrow}\right)=0$. By calculating the Chern numbers numerically, we obtain $C^{\uparrow}=-2$ and $C^{\downarrow}=2$ for our system, consistent with the second-order antiskyrmion configuration in Fig. 3-c). This is related to the presence of the $Gk_{\pm}^{2}$ terms in the off-diagonal matrix elements of the $\bm{k}\cdot\bm{p}$ Hamiltonian in Eq. (15). For $G=0$ (with the remaining parameters unchanged), the Chern numbers are given by $C^{\uparrow}=1$ and $C^{\downarrow}=-1$.
An important topological invariant used to characterize 2D TIs is the so-called spin Chern number [25, 26, 27, 28, 92, 93, 94, 95], defined as
@@
For our system, $C_{s}=-2$, which characterizes the presence of a TI phase. Note that, at zero temperature, the spin Hall conductivity within the bulk band gap is related to the spin Chern number via $\sigma_{s}=C_{s}(e/2\pi)$ [2]. Therefore, we predict that $\sigma_{s}=-2e/2\pi$ for Na2CdSn.
@@
As shown above, Na2CdSn hosts the TI and TCI phases simultaneously; thus, it can be classified as a DTI protected by TRS and mirror symmetry. Note that both topological invariants (spin and mirror Chern numbers) are equal to $-2$ (in sharp contrast with the spin and mirror Chern numbers $-1$ reported in Ref. [70]). According to the bulk-boundary correspondence [23, 24], the Na2CdSn triatomic layer in a finite geometry must have two pairs of counterpropagating helical edge states at each of its boundaries [85]. We explore this feature in the next section.
Unlike well-known conventional TIs [5, 10], the band inversion in Na2CdSn does not follow the closure of a trivial bulk band gap. Here, the topological gap arises from the SOC-induced lifting of a zone-centered nodal point [Figs. 1-d) and 1-e)]. The nodal point degeneracy is enforced by the 2D irrep $\Gamma_{6}$ of the low-energy conduction and valence bands. Once SOC is considered, the degeneracy is immediately lifted [Figs. 1-f) and 1-g)], resulting in a topologically nontrivial band gap characterized by the spin and mirror Chern numbers -2.
We note that the topology of Na2CdSn is fundamentally distinct from reported systems with similar crystal structures, such as Na2MnPb and $\alpha$-Bi. The former has been predicted to be a ferromagnetic DTI with $C_{s}=1$ and broken TRS [97], while the latter is predicted to have $C_{s}=2$ but no crystal-symmetry-related topological invariant [98]. Additionally, it is noteworthy that Ref. [85] reported a nontrivial spin Chern number $C_{s}=2$ in a generic model for 2D mirror-symmetric paramagnetic semiconductors. However, the mirror Chern number, proposed in Ref. [32], was not investigated in their study.
@@ V Topological edge states
Let us now investigate the existence of topological edge states predicted by the nontrivial topological invariants calculated in Sec. IV. In Sec. V.1, we analytically obtain the edge states for a semi-infinite plane geometry. In Sec. V.2, we perform a numerical study of the topological edge states in a ribbon geometry.
@@ V.1 Semi-infinite plane
In this geometry, any edge-localized eigenstate $\Psi_{k_{y}}(x,y)$ must satisfy the following boundary conditions:
@@
Equation (32) is the Born-von Karman boundary condition, with $L_{y}$ the length of the system in the $y$ direction.
Consider the Schrödinger equation for the spin-up block $h(k_{x}\rightarrow-i\partial_{x},k_{y})\psi(x,y)=E\psi(x,y)$, where $h(\bm{k})$ is defined by Eq. (25) and $\psi(x,y)$ is a particular two-component envelope function solution. The wave function $\psi(x,y)$ can be separated into a product of two factors, one corresponding to plane waves along the $y$ direction and the other corresponding to a nonperiodic function along $x$. This allows us to write the following ansatz [100]
$\psi_{\kappa_{x},k_{y}}(x,y)=\frac{\exp(ik_{y}y)}{\sqrt{L_{y}}}\begin{pmatrix}1\\ R_{\kappa_{x},k_{y}}\end{pmatrix}\exp(i\kappa_{x}x),$ (33)
@@
Substituting Eq. (33) into the Schrödinger equation, we obtain four modes given by $\kappa_{x}=(-1)^{j}\kappa_{\pm}$, where $j=\{0,1\}$ and
@@
where $d_{k_{y}\pm}=-M-\bar{E}+(B-D)(\kappa_{\pm}^{2}+k_{y}^{2})$. From now on, we use $R^{j}_{k_{y}\pm}\equiv R^{j}_{\pm}$. Using Eqs. (33), (34), and (38), we write the general solution as the linear combination of the four modes:
$\Psi_{k_{y}}(x,y)=$ $\frac{\exp(ik_{y}y)}{\sqrt{L_{y}}}$
×$\times\sum_{\begin{subarray}{c}j=0,1\\ l=+,-\end{subarray}}c_{j,l}\begin{pmatrix}1\\ R^{j}_{l}\end{pmatrix}\exp[i(-1)^{j}\kappa_{l}x].$ (39)
@@
Note that, in applying the boundary conditions in Eqs. (30)-(32), we have already neglected the exponential terms that do not result in square-integrable solutions, i.e., that do not yield a vector Hilbert space. Using Eqs. (39)-(42), we obtain
$\Psi_{k_{y}}(x,y)=$ $c_{0,+}\frac{\exp(ik_{y}y)}{\sqrt{L_{y}}}$
×$\times\begin{Bmatrix}\exp(i\kappa_{+}x)-\exp(i\kappa_{-}x)\\ R^{0}_{+}\left[\exp(i\kappa_{+}x)-\exp(i\kappa_{-}x)\right]\end{Bmatrix},$ (43)
where $c_{0,+}$ is determined by normalizing the wave function. Similarly, for $\text{Im}\left(\kappa_{-}\right)<0$, we have
@@
$\Psi_{k_{y}}(x,y)=$ $c_{0,+}\frac{\exp(ik_{y}y)}{\sqrt{L_{y}}}$
×$\times\begin{Bmatrix}\exp(i\kappa_{+}x)-\exp(-i\kappa_{-}x)\\ R^{0}_{+}\left[\exp(i\kappa_{+}x)-\exp(-i\kappa_{-}x)\right]\end{Bmatrix}.$ (47)
@@
In Fig. 4, we show the analytical dispersion relations of the spin-up (red curve) and down (blue) edge states [Eq. (48)] for the fitted parameters (Sec. III.3). The gray-shaded areas represent the continua of delocalized bulk states (bulk region), whose boundaries are indicated by solid black lines.
Figure 4: (Color online) Topological edge states hosted by the $\bm{k}\cdot\bm{p}$ model in a semi-infinite plane geometry, with parameters from the fitting for Na2CdSn. The red (blue) solid lines represent the analytical solutions for the dispersion relations of the spin-up (down) edge states [Eq. (48)], while the solid black lines delimit the bulk region, corresponding to the continua of delocalized bulk states (gray-shaded areas).
For each energy level inside the bulk band gap, there are two pairs of counterpropagating topological edge states of opposite spin. This is in agreement with the bulk-boundary correspondence theorem, considering the topological invariants calculated in Sec. IV. As it can be seen from Eq. (51), the quadratic dispersion relation of the topological edge states is a consequence of the nonzero parameter $G$.
Note that all edge states merge into the bulk region at $k_{y}\approx 0$. As $k_{y}$ moves away from zero, however, the energy of the edge states increases (in absolute value) slower than the energy of the bulk bands. Hence, here, the edge states never merge into the bulk region again. We emphasize that the effective model derived in Sec. III is only valid in the region close to $k_{y}=0$; thus, we should not expect it to capture any realistic merging points that might occur at larger values of $k_{y}$. The absence of merging points is also observed when we consider edge states arising from the BHZ Hamiltonian in the absence of $k^{2}$ diagonal terms [101].
@@ V.2 Ribbon geometry
In Fig. 5-a), we show the electronic band structure for a ribbon of width $W=100\,$nm, i.e., a wide ribbon (compared with the lattice parameter $a=4.978\,$Å). Each $(k_{y},E)$ point is colored according to its spin, with red (blue) representing spin-up (down) states. The energy bands forming a Mexican-hat-like structure constitute the bulk region. The energy gap $E_{g}\approx 285\,$meV between the lowest bulklike conduction and topmost bulklike valence bands is indicated in the figure.
@@
Figure 5: (Color online) Electronic states for Na2CdSn in a ribbon geometry. a) Electronic band structure showing spin-up (red) and spin-down (blue) states with a small spin splitting caused by the bulk inversion asymmetry (BIA) and $\bm{k}$-dependent spin-orbit coupling (SOC). The bulk region is determined by the Mexican-hat-like energy bands, with a bulk band gap of $E_{g}\approx 285$ meV. The quadratic energy bands correspond to quasidegenerate topological edge states, with the spin split states localized at opposite sides of the ribbon. The colored cross marks indicate the selected edge states for our wave function analysis, and the two black arrows denote the points of edge state crossings. b) Probability densities of the selected topological edge states as functions of $x$ (coordinate along the transversal direction). The color of the curves corresponds to the colors of the respective cross marks, while the solid (dashed) style represents spin-up (spin-down) states. c) Electronic band structure for Na2CdSn in a ribbon geometry of width $W=15\,$nm. The black arrows indicate the opening of an anticrossing hybridization gap due to the finite-size effect.
@@
The quadratic energy bands crossing the bulk band gap correspond to the topological edge states. All energy bands are quasidegenerate, including the edge states for which the spin splittings are $\sim 10^{-1}$ meV. Like the spin splitting of the energy bands in the infinite bulk system [Fig. 2-c)], the spin splitting of the edge states is not accompanied by a mixing of the spin components due to the block-diagonal structure of the effective Hamiltonian [Eq. (15)].
@@
The quasidegenerate topological edge states are localized at opposite sides of the strip. Note that, in the wide-ribbon limit, the energy dispersions of the edge states localized on the left (right) side can be mapped onto those of a vacuum/DTI (DTI/vacuum) semi-infinite system. In analogy to Sec. V.1, we employ the envelope function approach and solve the problem for the DTI/vacuum geometry. We use the analytical solutions for the two semi-infinite systems to derive an expression for the inter-edge splitting. Subtracting the dispersion relations for spin-up and spin-down states, we obtained that the spin splittings are given by $2b_{1,1}k_{y}$ and $2b_{2,1}k_{y}$ for the concave up and down bands, respectively.
@@
Notice that there is no structural inversion asymmetry in this case, which for the BHZ model is known to generate a Rashba term [103, 104, 105, 105, 106, 107] that lifts the degeneracy of the topological edge states [107]. For the ribbon geometry considered here, the spin splitting arises solely from the BIA. It can be shown that the inclusion of BIA in the BHZ Hamiltonian [100, 108] also leads to interedge spin splitting in 2D TIs, although in this case, the splitting is accompanied by a mixing of the spin components.
@@
In Fig. 5-b), we present the probability density of the edge states, indicated by the green and orange cross marks in Fig. 5-a), as functions of $x$ ($x=0$ and $x=W$ correspond to the ribbon extremities). Each cross mark represents two quasidegenerate states localized at opposite sides of the ribbon (the line colors match the colors of the corresponding cross mark). The dashed (solid) green line corresponds to a spin-down (spin-up) state localized at the left (right) side of the ribbon. Similarly, the solid (dashed) orange line corresponds to a spin-up (spin-down) state localized at the left (right) side of the ribbon. Note that the states shown by the green color are slightly more localized than the ones represented in orange  [^3] A zoom-in on the probability density distributions shown in Fig. 5-b) reveals that the green (orange) curves present a normal (oscillatory) exponential decay with $x$. The decaying behavior of a given edge state is dictated by the region where it is found in the electronic band structure. This effect is well-discussed for the BHZ model in Ref. [109]. In our model, however, such an analysis is much more complicated. Therefore, we plan to provide a more detailed investigation of this topic in a forthcoming work..
@@
So far, we have only investigated the edge states for negative crystal momentum ($k_{y}\approx-0.24\,$nm-1 and $k_{y}\approx-0.48\,$nm-1). Since our model is invariant under TRS, each topological edge state has a Kramers partner localized on the same side of the ribbon but with opposite momentum and spin. The yellow (black) cross mark in Fig. 5-a) indicates the two Kramers partners of the two states indicated by the green (orange) cross mark. We have verified that, in the wide-ribbon limit, the dispersion relations for the edge states localized at the left side of the ribbon agree exactly with those obtained for the semi-infinite plane geometry [see Eq. (48) and Fig. 4].
@@
For any energy level within the bulk band gap (e.g., at the Fermi level), there are four topological edge states at each side of the ribbon. On the left side, there are two spin-down (up) edge states propagating along the $+y$ ($-y$) direction, while on the right side, there are two spin-up (down) edge states propagating along the $+y$ ($-y$) direction.
The black arrows in Fig. 5-a) point to the two edge state crossings taking place around the same energy level $E\approx 0.123\,$eV. At each crossing point, there are four topological edge states. Let us focus on the edge state crossing around $k_{y}\approx-0.4\,$ nm-1 (left black arrow). Two of the four edge states live on the same side of the strip, have opposite spins, and propagate along opposite directions. Since they live in different spin subspaces, they are orthogonal to each other and do not hybridize. The counterpropagating states localized at opposite sides of the ribbon, on the other hand, belong to the same spin subspace, and therefore, can hybridize. Naturally, the hybridization requires a superposition of their wave functions, the so-called finite-size effect [110, 109]. In Fig. 5-c), we show the electronic band structure for Na2CdSn in a ribbon of width $W=15\,$nm. For this narrow ribbon, the edge states (with the same spin) living on opposite sides of the strip hybridize. The hybridization gap due to the anti-crossing of the edge state bands around $k_{y}\approx-0.4\,$nm-1 and $k_{y}\approx 0.4$ nm-1 are indicated by the two black arrows in the figure.